OpenStudy (anonymous):

How would you graph a circle with the equation x^2 + y^2 + 4x - 10y + 13 = 0?

OpenStudy (anonymous):

completing squares

OpenStudy (anonymous):

Ok so I have (x^2 - 4X ) + (y^2 - 10y )= -13? I am stuck on what to fill in the blanks in both parentheses with...

OpenStudy (anonymous):

To complete squares you wan the expression that squares to give the expressions you have in the brackets above so (x-2)^2 will give us the x^2 and 4x terms but also give us a 4 which we don't want so we have to take it out (subtract it). You do the same thing with the y^2. You end up with (x-2)^2 -4 +(y-5)^2 -25 = -13 and then bring all the integers over to rhs (x-2)^2 + (y-5)^2 = -13 +29 = 16 = 4^2 which is in standard form.